An object of 30 kg is moved with a velocity of 2 m/s on a horizontal smooth surface. What is the velocity of the block for 4 seconds if force of 40 N is applied on it in the direction of force?
1
Expert's answer
2020-04-20T15:11:14-0400

mass of body= 30 Kg, initial velocity of body= 2 m/s and Force acting on it=40 N for 4 sec

F=ma

40=30a

a=1.33 "\frac{m}{s^2}"

Now after 4 sec velocity become,


v=u+at

v= 2+1.33"\times4=7.32 m/s"

So, velocity after 4 sec becomes 7.32 m/s